Adam is x years old. His father is 25 years older than Adam. His brother is 3 years younger than Adam. The brothers' ages added together equal their father's age.
a) Write an equation using x.
b) Solve it to find how old Adam is.

Found 2 solutions by Theo, rothauserc:
Answer by Theo(13342)

let x = adam's age.
let x + 25 = his father's age.
let x - 3 = his brother's age.

the sum of the brothers' ages = the father's age.

the formula for that is x + x - 3 = x + 25.

combine like terms to get 2x - 3 = x + 25.

that would be the equation you are looking for using x.

subtract x from both sides of this equation and add 3 to both sides of this equation to get:

2x - x = 25 + 3.

combine like terms to get x = 28.

that's adam's age.
his father's age is 28 + 25 = 53.
his brother's age is 28 - 3 = 25.

the sum of adam's and his brother's age is 28 + 25 = 35.

that's equal to their father's age.

your solutions are:

a) Write an equation using x.

2x - 3 = x + 25

b) Solve it to find how old Adam is.

adam is 28 years old.

Answer by rothauserc(4718)

Adam is x years old
:
Father's age = x+25 years old
:
Adam's brother is x-3 years old
:
a) x +x-3 = x+25
:
b) x +x-3 = x+25
:
2x = x+28
:
x = 28
:
*****************************************
Adam is 28 years old
:
Adam's father is 28+25 = 53 years old
:
Adam's brother = 28-3 = 25 years old
*****************************************
: